The Method of Substitution
5. On Derwin farm there is a total of 50 chickens and cows.If there are 118 legs, how many chickens are there?
Can you please help with this question??

Answer by checkley77(12844)

CHICK+COWS=50
CHICKENS=50-COWS.
2CHICKENS+4COWS=118
2(50-COWS)+4COWS=118
100-2COWS+4COWS=118
2COWS=118-100
2COWS=18
C0WS=18/2
COWS=9 ANS.
CHICKENS+9=50
CHICKENS=50-9
CHICKENS=41 ANS.
PROOF:
2*41+4*9=118
82+36=118
118=118